OpenStudy (anonymous):

Which expression is equivalent to the given expression? 19x – 23x

OpenStudy (anonymous):

19x - 23x = - 4x

OpenStudy (anonymous):

Is that what you needed or are their answer choices ?

OpenStudy (anonymous):

the answer choices are: A. 19(x – 23) B. 19x(1 – 23x) C. x(19 – 23) D. x(23 – 19)

OpenStudy (anonymous):

so probably A?

OpenStudy (anonymous):

no...c is your answer A : 19(x - 23) = 19x - (19*23) -- not equivalent B : 19(1 - 23x) = 19 - (19*23x) -- not equivalent C : x(19 - 23) = 19x - 23x -- equivalent D: x( 23 - 19) = 23x - 19x -- not equivalent

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

anytime :)